Status Report

NASA Space Station On-Orbit Status 14 Nov 2003

By SpaceRef Editor
November 16, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 14 Nov 2003

All ISS systems continue to function nominally, except as noted previously or below.

The crew started the day with a review of the timeline for the upcoming (11/18) training/demo exercise of crewmembers transferring from the DC-1 “Pirs” docking compartment/airlock to the Soyuz TMA-3 Habitation Module (BO) while wearing inflated Orlan-M spacesuits. [In case of a lifeboat contingency during the Russian two-man EVA planned for January/February next year, the spacewalkers would have to ingress the Soyuz CRV (crew return vehicle) from the DC-1.  The suited inboard exercise will test and demonstrate the crewmembers’ capability to perform this Orlan transfer in case it should become necessary.]

As part of the suited drill, CDR/SO Michael Foale and FE Alexander Kaleri started preparing the DC-1 and Soyuz BO modules.  [In the DC-1, these preps included installation and location of specific equipment, with emphasis on freeing the pass way to the Soyuz.  In the BO, a payload frame was to be removed, equipment to be installed and located, and the hatch cover between BO and Soyuz Descent Module (SA) closed and open for test.]

Further preparations today included assembling and readying replaceable components and consumables (e.g., batteries, cooling water, oxygen tanks, feedwater filters, moisture collectors, LiOH canisters) for the two Orlan-M suits #14 (blue) for Foale and #23 (red) for Kaleri, along with their BRTA telemetry units.  Actual checkout and final preps of the Orlan-Ms are scheduled for Monday (11/17), followed by the exercise on 11/18.

Also in support of the suited transfer exercise, Kaleri disconnected the threaded QD (quick disconnect) screw clamps (which rigidize the mating surfaces) from the DC-1 side and inspected the hatch after their removal. 

The Flight Engineer completed the second session of the Russian MBI-8 “Profilaktika” (countermeasures) fitness program, consisting of the blood testing part, which determines lactate and creatine kinase levels in the blood with the AccuSport equipment, followed by the physical exercise session on the TVIS treadmill.  Mike Foale assisted as CMO (crew medical officer). [The TVIS test is identical to the MO-3 test performed on the treadmill in idling mode with free choice of speeds within a certain specified range.  In addition to the nominal test procedure, MBI-8/Part 3 calls for the use of the TEEM-100M gas analyzer during the test, the blood lactate measurements, and subjective evaluation of physical exertion levels during the test.  At the end of the creatine kinase tests, the results were logged and recorded in Kardiocassetta-2000 for subsequent downlink to the ground.]

Sasha performed the regular once-a-week maintenance reboot on the operational PCS laptops.  He also rebooted the OCA comm router laptop.

Kaleri called down the O2 partial pressure of the cabin air.  [“Ad hoc” O2 data are collected daily with the U.S. CSA-CP (compound specific analyzer-combustion products) for trending analyses by the ground.]

CDR Foale performed his first CBOSS-FDI (Cellular Biotechnology Operations Support System/Fluid Dynamics Investigation) session, first reviewing an OBT (onboard training) course for today’s task, then conducting a very simple procedure of labeling and making syringes.  The procedure was video taped for documentation, and the ground team stood by for support.

Sasha completed the regular weekly inspection of the active BRPK-1 air/liquid condensate separator system in the Service Module (SM).

Mike transferred data files from the physical exercise equipment to the MEC (medical equipment computer) via memory card and RED log entries, for downlink on OCA comm.  Afterwards, he also performed the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, then deleted them on the HRM.  

Foale also completed the regular status checkup of autonomous Lab payloads, while FE Kaleri conducted the daily routine maintenance of the SM’s SOZh life support system (including ASU toilet facilities) and updated the IMS “delta” file for updating the inventory databases.

A detailed picture of a UCD (urine collection device) of the HRF Renal Stone payload was uplinked for the crew to identify specific damage area(s) in onboard UCDs reported by them before.

The CDR performed the periodic inspection of the RED (resistive exercise device).

ISS flight attitude continues to be XPOP (Yaw 180.5 deg; Pitch 351.0 deg; Roll 0 deg), to be maintained until 11/27.

Today’s CEO (crew earth observation) targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, were West Africa Biomass Burning (Dynamic event:  Fires reported in the Sahel region, inland from the coast.  Looking left for this low sun angle [sunrise] opportunity),Angolan Biomass Burning(offshore air flow continues, with ground stations in the African subcontinent reporting reduced visibility under high atmospheric pressure.  Looking left towards Namibia for any haze masses moving into the Atlantic.  Oblique views, wide-angle lenses [or a series of pans along the horizon], and slight overexposure all help to image airborne particulates),Lower Amazon River Basin(nadir pass.  Crew was asked to shoot estuary islands and coastlines.  The former are highly mobile and the subject of new research),Fires, NE Brazil(Dynamic event:  Fires reported in arid NE Brazil),Stream patterns, Amazonia(sun-glint opportunity to capture river patterns of the Amazon’s major southern tributaries in SW Amazonia.  River positions and patterns vary ever further from mapped positions.  Looking left towards the glint disc.  Opportunity lasted for three minutes),Lake Eyre, Australia(Dynamic event.  Sun-glint pass: looking left about two degrees for this largest of Australia’s dry lakes, which may still have some water after recent flooding),Industrialized SE Africa smog (local stations reporting smog haze for the first time in many weeks: ideal sunset pass, looking left, over the Witwatersrand, the subcontinent’s economic hub), andPatagonian Glaciers(cloudfree slot should remain for one more day, allowing viewing of glaciers on both sides of the Andes; three passes/opportunities).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:39pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 18A.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 27.1; ppO2 (mmHg) — 157.8; ppCO2 (mmHg) — 3.9.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.7.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 23.0.
  • Node:  Pressure (mmHg) — 745.16; temperature (deg C) — 24.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 746.94; temperature (deg C) — 24.2; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 747.04; temperature (deg C) — 28.7; shell heater temp (deg C) — 25.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 25.3
  • PMA-2:  Shell heater temp (deg C) — 10.8

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack mode (suntracking). 
  • SM batteries:  Battery #8, formerly known as #7, is still in slot #8 for troubleshooting; all other batteries (7) are in “Partial Charge” mode (batteries #1 and #3 are degraded).  
  • FGB batteries:  Batteries #1 is off; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational (all lanes reintegrated 11/5).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; string #3 dropped out 10/22.

Propulsion System:

  • Total propellant load available:3734 kg (8232 lb) as of 11/6 [SM(755) + FGB(2627) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

SpaceRef staff editor.